OpenStudy (anonymous):

Evaluate the expression, using the given value of the variable. z ÷ 3 – 5z + 4(z – 2) when z = 12

OpenStudy (anonymous):

12= z

OpenStudy (anonymous):

so, 12/3 = 4

OpenStudy (anonymous):

4(12-2) = 40

OpenStudy (anonymous):

i know that its just i got -142 im really bad in math

OpenStudy (anonymous):

4-5(12)+40=?

OpenStudy (anonymous):

do multiplication first.

OpenStudy (anonymous):

It's fine :)

OpenStudy (anonymous):

i dont have my glasses on so i saw 12-2 as multiplaction

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

haha ;)

OpenStudy (anonymous):

No problem! :)